I need help with the following problem:
In the following sequence, what could be the next term? 1, 4, 27, 256, ... why is it the next term?
Any explanation would be helpful.
Thanks in advance.

Answer by ankor@dixie-net.com(22740)

How about:
1^1 = 1
2^2 = 4
3^3 = 27
4^4 = 256
5^5 = 3125
:
Starting with 1 the number is the result of raising consecutive numbers by the same value